# VFScale: Intrinsic Reasoning through Verifier-Free Test-time Scalable Diffusion Model

### v3

@@ A.3 Linear-regression algorithm
Given three points $(x_{1},y_{1})$, $(x_{2},y_{2})$, and $(x_{3},y_{3})$, we wish to fit a line of the form  [27]:
@@ A.4 Inference algorithm
Detailed algorithm of hMCTS of VFScale inference method is in Algorithm 1.
@@
11:        Expansion: Use Eq. 8 to generate child node state $\bm{x}_{t^{\prime}-1}^{(k)}$ for $\bm{x}_{t^{\prime}}$ and initialize $Q(\bm{x}_{t^{\prime}-1}^{(k)})=0$, $N(\bm{x}_{t^{\prime}-1}^{(k)})=0$ for $k=0,...,K-1$;
@@ Appendix B Details of experiments
B.1 Core Metric Definition
For Maze, successful solving means finding a continuous path from starting location to target location without any breaking or overlapping with the wall; for Sudoku, successful solving means filling all the missing numbers that exactly match the ground-truth, both of which are very stringent metrics.
B.2 Detais of Sudoku experiments
For Sudoku experiment, the dataset, model architecture, and training configurations are adopted from [5]. We mainly use solving success rate to evaluate different models. Model backbone and training configurations can be found in Fig. 5 and Table 6, respectively. All the exploration hyperparameters $c$ are set as 100 for Sudoku task.
@@ B.2 Detais of Sudoku experiments
Figure 5: The model architecture for VFScale on Sudoku task. The energy value is computed using the L2 norm of the final predicted output similar to [18], while the output is directly used as noise prediction for the diffusion baseline.
@@
B.3 Details of Maze experiments
The details of maze experiments and model backbone are provided in Table 7 and Fig. 6, respectively. The key metric, the maze-solving success rate is defined as the proportion of model-generated paths that have no breakpoints, do not overlap with walls, and begin and end at the start and target points, respectively. Maze datasets are generated by [24], and detailed hyperparameter configurations are in Table 7. All the exploration hyperparameters $c$ are set as 100 for Maze task.
@@ B.3 Details of Maze experiments
Figure 6: The model architecture for VFScale on Maze task. The energy value is computed using the L2 norm of the final predicted output similar to [18], while the output is directly used as noise prediction for the diffusion baseline.
@@ Appendix C Performance sensitivity to hyperparameters
In this subsection, we analyze the impact of several hyperparameters on the experimental results. As shown in Table 9, the influence of different noise scales on the performance of various methods is presented. The hMCTS denoising and BoN require a relatively larger noise scale to better expand the search space and improve final performance, while the diffusion model with naive inference performs best with a smaller noise scale. As demonstrated in Table 8 and Fig. 7, the effect of varying inner-loop optimization steps on the results is also analyzed. It can be observed that performance improves gradually with an increasing number of steps, and after 5 steps, the performance stabilizes and the improvement slows down. Therefore, we chose 5 inner-loop optimization steps for the Maze experiments in this paper.
@@ Appendix D Additional results
The parameter $t_{s}$ controls the proportion of the total inference budget allocated to MCTS denoising. When $t_{s}=9$, it means only MCTS denoising is used, while $t_{s}=0$ means only BoN is employed. For $90<t_{s}<$ 9, hMCTS denoising is applied. As shown in Table 10 and Fig. 8, there is a noticeable peak in model performance as $t_{s}$ varies.
@@
Table 11: Success rate of BoN for different training methods on Maze with grid size 15 and Sudoku harder dataset guided with ground truth accuracy. Untrained, BoN (gt) represents use ground truth to guide the BoN. Here, $L=N$. Bold font denotes the best model.
@@ Appendix E Extension of hMCTS to Conventional Diffusion Models
The inference method hMCTS of VFScale is not limited to energy-based diffusion models. With an external verifier, it can be applied to conventional models that predict noise directly. We validated this on the Maze task by replacing the learned energy with MSE as the reward, confirming hMCTS’s compatibility. Its gains over best-of-N search from the Table 13 below further demonstrate broad applicability across diffusion models.
Table 13: Success rate on Maze (grid size 15) using a conventional noise-predicting diffusion model (DDPM) guided by an external ground-truth verifier. We compare the performance of hMCTS with BoN under different compute budgets ($N$). Here, the reward is defined as the negative MSE to the ground-truth solution. Results validate the compatibility of hMCTS with conventional diffusion and its improved efficiency over best-of-$N$ sampling.
@@
Appendix F Broader Impacts

### v4

@@ A.3 Linear-regression algorithm
Given three points $(x_{1},y_{1})$, $(x_{2},y_{2})$, and $(x_{3},y_{3})$, we wish to fit a line of the form  Lane et al. (2003):
@@ A.4 Inference algorithm
The detailed algorithm of hMCTS of VFScale inference method is in Algorithm 1.
@@
11:   Expansion: Use Eq. 8 to generate child node state $\bm{x}_{t^{\prime}-1}^{(k)}$ for $\bm{x}_{t^{\prime}}$ and initialize $Q(\bm{x}_{t^{\prime}-1}^{(k)})=0$, $N(\bm{x}_{t^{\prime}-1}^{(k)})=0$ for k=0,,K1k=0,...,K-1;
@@ Appendix B Details of experiments
B.1 Overview
In Maze experiments, the datasets are generated by Ivanitskiy et al. (2023), and the diffusion model is mainly trained with Maze sizes of up to $6\times 6$ while tested on harder datasets with sizes significantly larger than $6\times 6$ (Fig. 1). For Sudoku experiments, the basic setting is adopted from Du et al. (2024) where the diffusion model is trained on SAT-Net dataset with 31 to 42 given entries Wang et al. (2019) and tested on the harder RRN dataset with 17 to 34 given entries Palm et al. (2018) with fewer given entries. All models and inference methods are evaluated with solving success rate. Here successful solving means the predicted solution exactly matches the ground-truth solution on all entries, a very stringent metric [^11] For example, a successful solving of a $15\times 15$ Maze needs to predict all $31\times 31=961$ entries correctly (for each grid cell, predict path/not path/wall), where the path length from start to target is on the order of $10^{2}$..
All training methods are compared with the original training pipeline in Du et al. (2024). LRNCL and KL represent the loss terms ${\mathcal{L}}_{\text{LRNCL}}$ and ${\mathcal{L}}_{\text{KL}}$, respectively in Eq. 6. VFScale tr. w/o LRNCL, VFScale tr. w/o KL, and VFScale tr. (ours) represent the three training methods of VFScale, the last being the full version. The MCTS denoising and hMCTS denoising are compared with BoN with the same computational budget. The experiment code can be found at the repo.
B.2 Core Metric Definition
For Maze, successful solving means finding a continuous path from starting location to target location without breaking or overlapping with the wall; for Sudoku, successful solving means filling all the missing numbers that exactly match the ground truth, both of which are very stringent metrics.
B.3 Details of Sudoku experiments
For Sudoku experiment, the dataset, model architecture, and training configurations are adopted from Du et al. (2024). We mainly use solving success rate to evaluate different models. The model backbone and training configurations can be found in Fig. 5 and Table 6, respectively. All exploration hyperparameters $c$ are set as 100 for the Sudoku task.
@@ B.3 Details of Sudoku experiments
Figure 5: The model architecture for VFScale on Sudoku task. The energy value is computed using the L2 norm of the final predicted output similar to Du et al. (2023), while the output is directly used as noise prediction for the diffusion baseline.
@@
B.4 Details of Maze experiments
The details of maze experiments and the model backbone are provided in Table 7 and Fig. 6, respectively. The key metric, the maze-solving success rate, is defined as the proportion of model-generated paths that have no breakpoints, do not overlap with walls, and begin and end at the start and target points, respectively. Maze datasets are generated by Ivanitskiy et al. (2023), and detailed hyperparameter configurations are in Table 7. All the exploration hyperparameters $c$ are set as 100 for Maze task.
@@ B.4 Details of Maze experiments
Figure 6: The model architecture for VFScale on Maze task. The energy value is computed using the L2 norm of the final predicted output similar to Du et al. (2023), while the output is directly used as noise prediction for the diffusion baseline.
@@ B.5 Compute Resources
B.6 Data Representation and Processing for Discrete Tasks
A key aspect of our "continuous reasoning formulation" is its application to discrete tasks like Maze (and Sudoku) via a discrete-continuous-discrete pipeline. This approach, which is common for applying generative models to discrete data (Du et al., 2024), allows the model to leverage gradient-based refinement in a continuous space. The process is as follows:
  • Discrete $\rightarrow$ Continuous Encoding: The discrete $H\times W$ Maze task is first encoded into a continuous tensor. We treat each grid cell as having one of $C$ states (e.g., "path," "wall," "empty") and use a one-hot encoding for each cell. This transforms the discrete $H\times W$ grid into a continuous $H\times W\times C$ tensor, $x_{0}$, which serves as the model’s ground truth.
  • Continuous Diffusion Process: Our energy function, $E_{\theta}$, and all search algorithms (BoN/hMCTS) operate entirely in this continuous $H\times W\times C$ space. All diffusion, noise prediction, and energy calculations are performed on these continuous-valued tensors.
  • Visualization (e.g., in Figure 2): To clarify our overview figure, the intermediate states shown (e.g., $x_{k}$) are not discrete paths. They are renderings of the continuous probability distribution (e.g., in the "path" channel) of the predicted $\hat{x}_{0}$ at an intermediate step. This visualizes the model’s evolving "belief" as it denoises from $t=T$ (pure noise) to $t=0$ (a confident solution).
  • Decoding to Discrete: To obtain the final discrete solution for evaluation, we apply a standard argmax operation along the channel dimension ($C$) of the final output tensor $\hat{x}_{0}$. This "collapses" the probability distribution back to the single most likely discrete state for each grid cell. This final discrete grid is then evaluated for success.
@@ Appendix C Performance sensitivity to hyperparameters
In this subsection, we analyze the impact of several hyperparameters on the experimental results. As shown in Table 9, the influence of different noise scales on the performance of various methods is presented. The hMCTS denoising and BoN require a relatively larger noise scale to better expand the search space and improve final performance, while the diffusion model with naive inference performs best with a smaller noise scale. As demonstrated in Table 8 and Fig. 7, the effect of varying inner-loop optimization steps on the results is also analyzed. It can be observed that performance improves gradually with an increasing number of steps, and after 5 steps, the performance stabilizes and the improvement slows down. Therefore, we chose 5 inner-loop optimization steps for the Maze experiments.
@@ Appendix D Additional results
D.1 The Challenge of External Verifiers and the Value of Intrinsic Energy
In the main text, we argue that leveraging the learned energy function as an intrinsic verifier is a cornerstone of our verifier-free, test-time scaling framework. This subsection provides a detailed comparative analysis to substantiate this claim, demonstrating two key points: (1) training a high-performing external verifier for complex reasoning tasks is nontrivial, and (2) the dense, continuous reward signal provided by our intrinsic energy is crucial for effective scaling and performs on par with a perfect oracle.
Intrinsic Verifier vs. Learned External Verifier: To assess the difficulty and effectiveness of using a separately trained model as a verifier, we constructed a learnable external verifier. Specifically, we trained a classifier to predict the element-wise accuracy of a given Maze solution. While this classifier achieved a high correlation of 0.99 with the true ground truth scores, its performance as a guide for test-time scaling was substantially inferior to our intrinsic energy verifier.
As shown in Table 10, when using Best-of-N (BoN) sampling, the external verifier leads to a performance drop of up to 30% compared to using the intrinsic energy function. We attribute this significant gap to the external verifier’s inability to distinguish between the subtle yet critical differences among high-quality candidate solutions generated during the scaling process, leading to suboptimal selections. This experiment highlights that even a verifier with high statistical correlation may fail to provide the precise guidance needed for complex reasoning.
Table 10: Comparison of test-time scaling performance on the $15\times 15$ Maze task using our intrinsic energy verifier versus a trained external verifier. Both methods use the same set of generated samples under the Best-of-N (BoN) framework. Success rates are reported for different compute budgets (NN).
Method N=1 N=11 N=21 N=41 N=81
VFScale tr. (ours), BoN, Intrinsic Verifier 0.2656 0.5859 0.6406 0.6875 0.6562
VFScale tr. (ours), BoN, External Verifier 0.2656 0.3516 0.3594 0.3438 0.3203
The Importance of a Dense Reward Signal:To further locate the advantage of our intrinsic verifier, we compared its performance against two perfect ground-truth (GT) oracles:
  • A sparse GT 0/1 verifier, which provides a binary success/failure signal only.
  • A continuous GT score verifier, which provides a precise continuous quality score for any given sample (a perfect, dense reward).
The results in Table 11 are highly informative. First, the sparse 0/1 verifier performs significantly worse (over a 10% gap at $N=81$) than our dense, energy-guided method. This starkly illustrates the advantage of continuous guidance; a dense reward signal allows the search algorithm (hMCTS) to make more informed decisions at each step, rather than relying on a simple binary outcome.
Crucially, the performance of our intrinsic verifier is nearly identical to that of the perfect continuous GT score verifier. This result strongly validates the effectiveness of our training objectives (especially LRNCL) in shaping a high-quality energy landscape that accurately reflects the true solution quality. In essence, our framework successfully learns an intrinsic verifier that functions as a near-perfect, dense reward oracle, unlocking the full potential of test-time scaling without external supervision.
Table 11: Performance comparison on the $15\times 15$ Maze task between our energy-guided intrinsic verifier and two perfect ground-truth (GT) oracles: a sparse binary verifier and a continuous score verifier. All methods use the hMCTS search algorithm.
Method N=1 N=11 N=21 N=41 N=81
VFScale (ours), hMCTS, Energy Guided (Intrinsic) 0.2656 0.6406 0.7266 0.7969 0.8203
VFScale (ours), hMCTS, GT 0/1 Guided (Sparse) 0.2656 0.5469 0.5938 0.6562 0.6719
VFScale (ours), hMCTS, GT Score Guided (Dense) 0.2656 0.6328 0.7266 0.7734 0.8359
D.2 Ablation Study on Monotonic Regression Constraints for MRNCL
To validate the effectiveness of the linear constraint and analyze the impact of different energy-L2 distance relationships within our Monotonic Regression Negative Contrastive Learning (MRNCL) framework, we conducted a comprehensive ablation study on the Maze task. We compare our proposed LRNCL (Linear) against several variants:
  • QRNCL (Quadratic Regression NCL): This method fits a quadratic function, encouraging the energy-L2 distance of samples to lie within its monotonically increasing region.
  • RBNCL (Rank-based NCL): This uses the rank correlation between sample energies and their L2 distances as a loss, without assuming a specific functional form for the relationship.
  • FTNCL (Fixed-transformation NCL): This requires the learned energy function to approximate a pre-defined linear function with fixed, non-adaptive parameters.
Table 12: Success rates on the $15\times 15$ Maze task for different MRNCL variants. While QRNCL achieves the best performance for single-shot generation (N=1), LRNCL demonstrates superior test-time scalability as the compute budget ($N$) increases, achieving the highest success rate at N=81.
Training Method N=1 N=11 N=21 N=41 N=81
Original, BoN, Energy Guided 0.0625 0.0469 0.0547 0.0781 0.1016
+LRNCL (ours) 0.2812 0.4688 0.4922 0.5547 0.5859
+QRNCL 0.3516 0.3672 0.3906 0.3984 0.3828
+RBNCL 0.1875 0.4219 0.4766 0.5234 0.5781
+FTNCL 0.3125 0.4141 0.4844 0.4922 0.5156
The scaling results in Table 12 show that all MRNCL variants significantly improve performance over the baseline. Notably, an interesting trade-off emerges: QRNCL yields the largest performance gain without scaling (N=1), but offers the weakest benefit for test-time scalability, with the performance even degrading at N=81. In contrast, RBNCL exhibits strong scalability, surpassed only by LRNCL, which provides the largest and most consistent scalability gain.
These findings collectively demonstrate that our proposed LRNCL, chosen for its simplicity, is highly effective at calibrating the energy landscape to improve a model’s test-time scalability, which is the primary goal of this work.
D.3 Analysis on switch-over step $t_{s}$ of hMCTS
@@
The parameter $t_{s}$ controls the proportion of the total inference budget allocated to MCTS denoising. When $t_{s}=9$, it means only MCTS denoising is used, while $t_{s}=0$ means only BoN is used. For $0<t_{s}<9$ (maximum denosing step is 10.), hMCTS denoising is applied. As shown in Table 14 and Fig. 8, there is a noticeable peak in model performance as $t_{s}$ varies.
While a manually tuned $t_{s}$ is effective, we also developed an adaptive mechanism to set the switch-over step on a per-sample basis, eliminating the need for hyperparameter tuning. This method achieves a more dynamic exploration-exploitation balance by using Best-of-N (BoN) for early-stage global exploration and switching to MCTS for later-stage local exploitation based on the state of the search.
Specifically, the mechanism utilizes the energy variance across the different search branches. The switch from BoN to MCTS occurs when this energy variance exceeds a preset threshold, creating a closed-loop system that dynamically balances the two strategies. As shown in Table 13, this adaptive approach shows a slight advantage over a manually tuned $t_{s}$, particularly with smaller inference budgets (e.g., at N=11 and N=21), demonstrating its utility and robustness.
Table 13: Success rate on the $15\times 15$ Maze task comparing a manually tuned, fixed switch-over step ($t_{s}$) with our proposed adaptive $t_{s}$ mechanism. The adaptive method shows a notable advantage at smaller compute budgets.
Switch-Over Method N=11 N=21 N=41 N=81
VFScale tr. (ours), hMCTS, Tuned $t_{s}$ 0.6406 0.7266 0.7969 0.8203
VFScale tr. (ours), hMCTS, Adaptive $t_{s}$ 0.7031 0.7734 0.7891 0.8203
@@
Table 15: Success rate of BoN for different training methods on Maze with grid size 15 and Sudoku harder dataset guided with ground truth accuracy. Untrained, BoN (gt) represents using ground truth to guide the BoN. Here, $L=N$. Bold font denotes the best model.
@@
D.4 Impact of Negative Sample Generation Strategy
To analyze the sensitivity of our framework to the negative sampling strategy, we compare the permutation-based approach used in our main experiments against a more standard Gaussian noise perturbation method. In this alternative, negative samples are generated by adding scaled Gaussian noise to the positive samples.
The results presented in Table 17 lead to two key findings. First, the LRNCL framework robustly improves performance and unlocks test-time scalability regardless of the specific negative sample generation strategy employed, demonstrating its general applicability. Both perturbation methods significantly outperform the original baseline.
Second, the choice of strategy is not without impact. While the Gaussian noise approach provides a strong performance boost, the permutation-based method yields superior test-time scalability, with its advantage growing as the compute budget ($N$) increases. This suggests that structured, problem-aware perturbations (like permutation noise for grid-based tasks) can create more informative negative samples for training the energy landscape compared to unstructured noise.
Table 17: Success rate on the $15\times 15$ Maze task comparing different negative sample generation strategies for LRNCL. While both strategies significantly outperform the baseline, the Permutation-based method shows better scalability.
Training Method N=1 N=11 N=21 N=41 N=81
Original, BoN, Energy Guided 0.0625 0.0469 0.0547 0.0781 0.1016
Permutation-based Perturb (ours) 0.2812 0.4688 0.4922 0.5547 0.5859
Gaussian Noise Perturb 0.3125 0.3750 0.3906 0.4062 0.4141
D.5 Stability Analysis and Multi-Seed Results
To validate the stability of our framework, we computed the standard deviation over 10 runs with different random seeds for our key experiments on the $15\times 15$ Maze task. The results, presented in Table 18, summarize the performance (mean $\pm$ std) for the primary methods as the compute budget (N) increases.
Table 18: Mean and standard deviation of success rates on the $15\times 15$ Maze task over 10 random seeds.
Methods N=11 N=21 N=41 N=81
Original, BoN 0.0539 $\pm$ 0.0055 0.0594 $\pm$ 0.0087 0.0594 $\pm$ 0.0063 0.0609 $\pm$ 0.0058
VFScale tr. (ours), BoN 0.6203 $\pm$ 0.0230 0.6562 $\pm$ 0.0148 0.6680 $\pm$ 0.0176 0.6836 $\pm$ 0.0165
VFScale tr. (ours), hMCTS (ours) 0.6727 $\pm$ 0.0230 0.7336 $\pm$ 0.0180 0.7945 $\pm$ 0.0086 0.8422 $\pm$ 0.0147
The data confirms that our VFScale framework is not only superior in performance but also stable, as indicated by the tight standard deviations. Both our training (VFScale tr.) and inference (hMCTS) methods show significant and consistent gains over the baseline, reinforcing the robustness of our approach.
D.6 From Local Constraints to Global Ranking Alignment
A natural question arises regarding our training objective: since MRNCL enforces only local monotonicity (between a positive sample and its perturbed negatives), how does this translate to a globally aligned energy landscape? Our hypothesis is that by repeatedly enforcing this simple, local, linear consistency across billions of samples, batches, and diffusion timesteps during training, we implicitly shape a globally structured and reliable energy landscape. We provide both direct and indirect evidence to support this.
Direct Evidence: Global Consistency Metrics. First, our Performance-Energy Consistency (PEC) metric (defined in Appendix A) serves as a measure of global rank-order consistency across a batch of samples. As shown in Table 4, our training improves this global consistency metric from $\sim 73\%$ (Original) to $\sim 84\%$ (VFScale).
To further rigorously quantify this, we conducted a new experiment measuring the Kendall-$\tau$ rank correlation—a standard metric for global ranking alignment—between the energy values and the true solution quality across various denoising timesteps. As presented in Table 19, our full method (“Learned energy”) achieves a significantly higher and more positive correlation compared to both the external verifier and the baseline energy model without MRNCL. Notably, in the critical intermediate stages of denoising ($t=0$ to $t=7$), our method maintains a strong positive correlation (>0.4>0.4), indicating that the energy landscape reliably guides the search towards better solutions globally.
Table 19: Kendall-$\tau$ rank correlation (mean and std) between the verifier score/energy and the ground-truth quality across denoising steps $t$. Our intrinsic energy verifier demonstrates superior global alignment compared to an external verifier and the baseline (w/o MRNCL).
Denoising Step $t$ Learned External Verifier Intrinsic Energy (w/o MRNCL) Intrinsic Energy (Ours)
0 -0.0978 (0.1750) 0.2300 (0.0854) 0.4760 (0.1492)
1 -0.0976 (0.1750) 0.2300 (0.0854) 0.4744 (0.1490)
2 -0.0980 (0.1756) 0.2308 (0.0854) 0.4760 (0.1492)
3 -0.0910 (0.1834) 0.2290 (0.0838) 0.4642 (0.1362)
4 -0.0824 (0.1678) 0.2368 (0.0800) 0.4094 (0.1538)
5 -0.1292 (0.1718) 0.2506 (0.0594) 0.4414 (0.1300)
6 -0.1318 (0.1748) 0.2638 (0.0766) 0.3978 (0.2466)
7 -0.3618 (0.1446) 0.2262 (0.1194) 0.4508 (0.2024)
8 -0.0316 (0.2854) 0.3214 (0.2138) -0.1570 (0.1852)
9 -0.2686 (0.0322) 0.3712 (0.0520) -0.3094 (0.0368)
Indirect Evidence: Search Success. The most conclusive proof of global alignment is the final outcome. The fact that our framework successfully scales—improving the $15\times 15$ Maze success rate from a baseline of 6.25% to 88.28% (Table 5)—serves as strong empirical evidence that a global change to the energy landscape has occurred. Such a dramatic improvement in global search performance would be unattainable if the energy landscape remained locally fragmented and globally unaligned.
@@ Appendix E Extension of hMCTS to Conventional Diffusion Models
The inference method hMCTS of VFScale is not limited to energy-based diffusion models. With an external verifier, it can be applied to conventional models that predict noise directly. We validated this in the Maze task by replacing the learned energy with MSE as the reward, confirming hMCTS’s compatibility. Its gains over best-of-N search from the Table 20 below further demonstrate broad applicability across diffusion models.
Table 20: Success rate on Maze (grid size 15) using a conventional noise-predicting diffusion model (DDPM) guided by an external ground-truth verifier. We compare the performance of hMCTS with BoN under different compute budgets (NN). Here, the reward is defined as the negative MSE to the ground-truth solution. Results validate the compatibility of hMCTS with conventional diffusion and its improved efficiency over best-of-NN sampling.
@@
Appendix F Computational Cost and Fair Comparison
This section provides a transparent analysis of the computational costs associated with the VFScale framework, covering both training and inference. We also detail our methodology for ensuring a fair comparison between different inference-time search strategies. All benchmarks were run on a machine with one 80GB VRAM GPU and 16 CPU cores.
F.1 Training Cost
By design, VFScale tackles a more complex training objective than the baseline to shape a search-friendly energy landscape. This increased complexity results in a higher demand for training time and memory, as detailed in Table 21 and Table 22.
While the LRNCL and KL losses increase the training cost, this upfront investment is a crucial trade-off. For the original method, performance plateaus once the loss converges, regardless of additional training. In contrast, our approach enables the base model’s performance to continuously improve with a larger training budget, which in turn unlocks significantly enhanced test-time scalability.
Table 21: Training time in hours for Sudoku and Maze task. The VFScale training objectives require more time but lead to models with superior scalability.
Method 100,000 Steps (hours) 200,000 Steps (hours)
Original 2 4
+LRNCL 4 5
+KL 3 6
+LRNCL+KL 6 7
Table 22: Peak GPU memory usage (VRAM in GB) during training with a batch size of 64.
Method Sudoku (GB) Maze (GB)
Original 3.28 3.40
+LRNCL 4.56 5.86
+KL 4.44 21.21
+LRNCL+KL 5.57 23.05
F.2 Inference Cost and Fair Comparison
To ensure a fair and meaningful comparison between inference methods (BoN, MCTS, and our hMCTS), we allocate the same Number of Function Evaluations (NFE) to each method for a given problem instance. Since the forward pass through the model is the primary computational bottleneck, controlling for NFE allows us to isolate the efficiency of the search strategy itself.
Table 23 reports the wall-clock inference time. As expected, hMCTS exhibits a modest overhead (up to  31% longer than BoN at N=81). This is a known trade-off, as the sequential nature of tree expansion in MCTS-based methods is inherently less parallelizable than the embarrassingly parallel BoN approach. However, hMCTS is still significantly more efficient than a pure MCTS implementation.
Crucially, our central claim holds: for an identical computational budget (NFE), hMCTS delivers superior performance over both BoN and MCTS. The difference in wall-clock time stems from the search strategy’s implementation rather than from the core computational complexity. We view this as a manageable engineering trade-off for the substantial performance gains achieved.
Table 23: Wall-clock inference time per sample (seconds) on the Maze task. For the same NFE, hMCTS provides superior results, justifying its modest time overhead compared to BoN.
Method N=1 N=11 N=21 N=41 N=81
BoN 2.17 2.80 5.63 11.01 20.16
hMCTS (ours) 1.02 3.28 6.56 12.66 26.48
MCTS 1.44 7.59 14.03 28.56 57.66
Appendix G Broader Impacts